# Wave-RVFL: A Randomized Neural Network Based on Wave Loss Function

### v1

@@ Abstract
The random vector functional link (RVFL) network is well-regarded for its strong generalization capabilities in the field of machine learning. However, its inherent dependencies on the square loss function make it susceptible to noise and outliers. Furthermore, the calculation of RVFL’s unknown parameters necessitates matrix inversion of the entire training sample, which constrains its scalability. To address these challenges, we propose the Wave-RVFL, an RVFL model incorporating the wave loss function. We formulate and solve the proposed optimization problem of the Wave-RVFL using the adaptive moment estimation (Adam) algorithm in a way that successfully eliminates the requirement for matrix inversion and significantly enhances scalability. The Wave-RVFL exhibits robustness against noise and outliers by preventing over-penalization of deviations, thereby maintaining a balanced approach to managing noise and outliers. The proposed Wave-RVFL model is evaluated on multiple UCI datasets, both with and without the addition of noise and outliers, across various domains and sizes. Empirical results affirm the superior performance and robustness of the Wave-RVFL compared to baseline models, establishing it as a highly effective and scalable classification solution.
@@ 1 Introduction
Artificial neural networks (ANNs) are machine learning models designed to mimic the structure and function of the human brain’s neural system. In ANNs, nodes, or “neurons”, are connected in layers that collaborate to process, analyze, and transmit information, enabling the network to make decisions. ANNs have shown success in various fields, including stock market prediction [12], diagnosis of Alzheimer’s disease [22, 23, 20] and significant memory concern [19], solving differential equations [3], and among others.
@@
Despite their advantages, ANN models face several challenges, such as slow convergence, difficulties with local minima, and sensitivity to learning rates [17]. To address these issues, randomized neural networks (RNNs) [21], such as random vector functional link (RVFL) neural networks [16], have been introduced. The RVFL model features a single hidden layer with weights from the input to the hidden layer that are randomly generated and fixed during training. It also includes direct connections from the input layer to the output layer, which serve as a built-in regularization mechanism, enhancing the RVFL’s generalization ability [26]. The training process focuses solely on determining the weights of the output layer, which can be efficiently accomplished using closed-form or iterative methods.
@@
Several enhanced versions of the standard RVFL model have been developed to improve its generalization performance, making it more robust and effective for practical applications [15]. The traditional RVFL model assigns equal weight to each sample, which makes it susceptible to noise and outliers. To address this issue, a modified version called intuitionistic fuzzy RVFL (IFRVFL) was proposed in [14], which uses fuzzy membership and nonmembership functions to assign an intuitionistic fuzzy score to each sample. In the standard RVFL, original features are transformed into randomized features, leading to potential instability. To counter this, Zhang et al. [27] incorporated a sparse autoencoder with $l1$-norm regularization into the RVFL, resulting in the SP-RVFL model. This model mitigates instability caused by randomization and enhances the learning of network parameters compared to the traditional RVFL. Li et al. [11] proposed another variant, the discriminative manifold RVFL (DMRVFL), which employs manifold learning techniques. DMRVFL replaces the rigid one-hot label matrix with a flexible soft label matrix to better utilize intraclass discriminative information and increase the distance between interclass samples. Recently, In [13], authors developed the graph embedded intuitionistic fuzzy weighted RVFL (GE-IFWRVFL), which integrates subspace learning criteria within a graph embedding framework. This approach aims to improve the RVFL’s performance by leveraging the graphical information of the dataset into the learning process. Additionally, to address class imbalance issues associated with RVFL, Ganaie et al. [6] proposed the graph embedded intuitionistic fuzzy RVFL for class imbalance learning (GE-IFRVFL-CIL). This model incorporates a weighting scheme based on the class imbalance ratio to manage class imbalance effectively. To incorporate human-like decision-making capabilities, Sajid et al. [18] introduced the neuro-fuzzy RVFL, which operates on the IF-THEN logic principle, thereby enhancing the interpretability of the RVFL model.
@@
Inspired by these promising properties of the wave loss function and motivated by the potential answers to these questions, we propose the Wave-RVFL, an RVFL model based on the wave loss function. In developing Wave-RVFL, we aimed to preserve the inherent optimization formulation of the RVFL as much as possible, simply replacing the square loss function with the wave loss function. The proposed optimization problem of Wave-RVFL is solved using the adaptive moment estimation (Adam) algorithm [10].
@@ 2.2 Random Vector Functional Link (RVFL) network
The optimal solution of Eq. (3) is defined as follows:
@@ 3 Proposed Work
In this section, we propose a novel model called the random vector functional link network with wave loss function (Wave-RVFL) by employing the wave loss function to give a penalty while creating the classifier. We further provide the mathematical formulation, its solution using Adam algorithms and the proposed algorithm. Wave-RVFL capitalizes on the asymmetry inherent in the wave loss function to dynamically apply penalties at the instance level for misclassified samples. The proposed Wave-RVFL demonstrates robustness against noise and outliers by preventing excessive penalization of deviations, thereby ensuring a balanced treatment of noise and outliers.
@@ 3.1 Wave loss function
    $\mathcal{L}_{0-1}(v)=\left\{\begin{array}[]{lr}0,&\text{if}~{}v\leq 0,\\ \frac{1}{\eta},&\text{if}~{}v>0.\end{array}\right.$ (10)
@@
3.2 Optimization Problem: Random Vector Functional Link Network with wave loss (Wave-RVFL)
@@
$s.t.~{}~{}z_{i}\beta=[x_{i},h(x_{i})]\beta-Y_{i}=\xi_{i},~{}\forall i\in\{1,2,\ldots,n\},$ (11)
@@
Optimizing the Wave-RVFL model presents challenges due to the non-convexity of the loss function. Nevertheless, by leveraging the inherent smoothness of Wave-RVFL, gradient-based algorithms can be effectively employed for model optimization.
@@ 3.3 Soultion: Wave-RVFL
At each iteration, $t$, $s$ samples are selected randomly and used to find the gradient and the following steps. Taking the gradient with respect to $\beta$, we obtain:
@@
Here, $\lambda_{1}$ and $\lambda_{2}$ represent the decay rates for the first and second-moment estimates, commonly set to $0.9$ and $0.999$, respectively. Subsequently, Next, we compute the bias-corrected first and second moment estimates using the following expression:
@@
Here, $\epsilon$ is a small constant, and $\alpha$ represents the learning rate.
@@
Input: Let $\{x_{i}\}_{i=1}^{n}$ be the input training dataset, $Y_{i}$ be the target output and $\mathcal{C}$, $\eta$, and $\gamma$ are the trade-off parameters, respectively.
@@ 4.1 Compared Models, Datasets and Experimental Setup
To evaluate the efficacy of the proposed Wave-RVFL models, we employ $23$ benchmark datasets from UCI [5] repository from various domians and sizes. For example, the number of samples in the “fertility” dataset is $100$ and the number of samples in the “connect_4” dataset is $67,557$. The experimental setup is discussed in Section S.II of the supplementary material.
@@ 4.2 Experimental Results on UCI Dataset
Considering the above discussion based on accuracy, rank, and statistical test, we can conclude that the proposed Wave-RVFL model showcases superior and robust performance against the baseline models.
@@ 4.3 Robustness Evaluation of the Proposed Wave-RVFL Model on Datasets with Outliers
    “magic” Dataset: The Wave-RVFL model led the performance table in $4$ out of $4$ scenarios for the magic dataset. It recorded the highest average accuracy at $63.3254\%$, showcasing its effectiveness under varying levels of label noise.
@@
These findings illustrate that the proposed Wave-RVFL model displays exceptional robustness against outliers, outperforming baseline models even in challenging conditions.
@@ 4.4 Robustness Evaluation of the Proposed Wave-RVFL Model on Datasets with Gaussian Noise
To evaluate the robustness of the proposed Wave-RVFL model against outliers, we introduced Gaussian noise into the magic and blood datasets at varying levels of $5\%$, $10\%$, $15\%$, and $20\%$. This approach provided a thorough comparison of the model’s performance under adverse conditions against the RVFL, as shown in Table 3.
@@
    Overall Performance: The Wave-RVFL model’s overall average accuracy was $72.0127\%$, approximately $10\%$ higher than the RVFL model. Furthermore, it attained the lowest overall average rank of $1.25$, underscoring its superior performance among all tested models.
@@
Figure 3: Effect of hyperparameters on the performance of the proposed Wave-RVFL model.

### v2

@@ Abstract
The random vector functional link (RVFL) network is well-regarded for its strong generalization capabilities in the field of machine learning. However, its inherent dependencies on the square loss function make it susceptible to noise and outliers. Furthermore, the calculation of RVFL’s unknown parameters necessitates matrix inversion of the entire training sample, which constrains its scalability. To address these challenges, we propose the Wave-RVFL, an RVFL model incorporating the wave loss function. We formulate and solve the proposed optimization problem of the Wave-RVFL using the adaptive moment estimation (Adam) algorithm in a way that successfully eliminates the requirement for matrix inversion and significantly enhances scalability. The Wave-RVFL exhibits robustness against noise and outliers by preventing over-penalization of deviations, thereby maintaining a balanced approach to managing noise and outliers. The proposed Wave-RVFL model is evaluated on multiple UCI datasets, both with and without the addition of noise and outliers, across various domains and sizes. Empirical results affirm the superior performance and robustness of the Wave-RVFL compared to baseline models, establishing it as a highly effective and scalable classification solution. The source codes and the Supplementary Material are available at https://github.com/mtanveer1/Wave-RVFL
@@ 1 Introduction
Artificial neural networks (ANNs) are machine learning models designed to mimic the structure and function of the human brain’s neural system. In ANNs, nodes, or “neurons”, are connected in layers that collaborate to process, analyze, and transmit information, enabling the network to make decisions. ANNs have shown success in various fields, including stock market prediction [12], diagnosis of Alzheimer’s disease [25, 26, 23], significant memory concern [21], speech enhancement [9], DNA-binding proteins [18] and among others.
@@
Despite their advantages, ANN models face several challenges, such as slow convergence, difficulties with local minima, and sensitivity to learning rates [19, 22]. To address these issues, randomized neural networks (RNNs) [24], such as random vector functional link (RVFL) neural networks [17], have been introduced. The RVFL model features a single hidden layer with weights from the input to the hidden layer that are randomly generated and fixed during training. It also includes direct connections from the input layer to the output layer, which serve as a built-in regularization mechanism, enhancing the RVFL’s generalization ability [29]. The training process focuses solely on determining the weights of the output layer, which can be efficiently accomplished using closed-form or iterative methods.
@@
Several enhanced versions of the standard RVFL model have been developed to improve its generalization performance, making it more robust and effective for practical applications [14]. The traditional RVFL model assigns equal weight to each sample, which makes it susceptible to noise and outliers. To address this issue, a modified version called intuitionistic fuzzy RVFL (IFRVFL) was proposed in [15], which uses fuzzy membership and nonmembership functions to assign an intuitionistic fuzzy score to each sample. In the standard RVFL, original features are transformed into randomized features, leading to potential instability. To counter this, Zhang et al. [30] incorporated a sparse autoencoder with $l1$-norm regularization into the RVFL, resulting in the SP-RVFL model. This model mitigates instability caused by randomization and enhances the learning of network parameters compared to the traditional RVFL. Li et al. [11] proposed another variant, the discriminative manifold RVFL (DMRVFL), which employs manifold learning techniques. DMRVFL replaces the rigid one-hot label matrix with a flexible soft label matrix to better utilize intraclass discriminative information and increase the distance between interclass samples. Recently, In [13], authors developed the graph embedded intuitionistic fuzzy weighted RVFL (GE-IFWRVFL), which integrates subspace learning criteria within a graph embedding framework. This approach aims to improve the RVFL’s performance by leveraging the graphical information of the dataset into the learning process. Additionally, to address class imbalance issues associated with RVFL, Ganaie et al. [5] proposed the graph embedded intuitionistic fuzzy RVFL for class imbalance learning (GE-IFRVFL-CIL). This model incorporates a weighting scheme based on the class imbalance ratio to manage class imbalance effectively. To incorporate human-like decision-making capabilities, Sajid et al. [20] introduced the neuro-fuzzy RVFL, which operates on the IF-THEN logic principle, thereby enhancing the interpretability of the RVFL model. Further, an efficient angle-based twin RVFL (ATRVFL) is proposed in [16] to effectively tackle binary classification problems.
@@
Inspired by these promising properties of the wave loss function and motivated by the potential answers to these questions, we propose the Wave-RVFL model, an RVFL model based on the wave loss function. In developing Wave-RVFL, we aimed to preserve the inherent optimization formulation of the RVFL as much as possible, simply replacing the square loss function with the wave loss function. The proposed optimization problem of Wave-RVFL is solved using the adaptive moment estimation (Adam) algorithm [10].
@@ 2.2 Random Vector Functional Link (RVFL) Network
The optimal solution of Eq. (3) can be calculated as follows:
@@ 3 Proposed Work
In this section, we propose a novel Wave-RVFL model, i.e., the random vector functional link network based on the wave loss function. We further provide the mathematical formulation, its solution using Adam algorithms and the proposed algorithm. Wave-RVFL capitalizes on the asymmetry inherent in the wave loss function to dynamically apply penalties at the instance level for misclassified samples. The proposed Wave-RVFL demonstrates robustness against noise and outliers by preventing excessive penalization of deviations, thereby ensuring a balanced treatment of noise and outliers.
@@ 3.1 Wave Loss Function
    $\mathcal{L}_{0-1}(v)=\left\{\begin{array}[]{lr}0,&\text{if}\leavevmode\nobreak\ v\leq 0,\\ \frac{1}{\eta},&\text{if}\leavevmode\nobreak\ v>0.\end{array}\right.$ (10)
@@
3.2 Optimization Problem: Random Vector Functional Link Network with Wave Loss Function (Wave-RVFL)
@@
$s.t.\leavevmode\nobreak\ \leavevmode\nobreak\ z_{i}\beta=[x_{i},h(x_{i})]\beta-Y_{i}=\xi_{i},\leavevmode\nobreak\ \forall i\in\{1,2,\ldots,n\},$ (11)
@@
Optimizing the Wave-RVFL model presents challenges due to the non-convexity of the loss function. Nevertheless, by leveraging the inherent smoothness of Wave-RVFL, gradient-based algorithms can be employed for model optimization.
@@ 3.3 Soultion: Wave-RVFL
At $t^{th}$ iteration (maximum number of iterations is denoted as “$Itr$”), $s$ samples are selected randomly and used to find the gradient and the following steps. Taking the gradient with respect to $\beta$, we obtain:
@@
Here, $\lambda_{1}$ and $\lambda_{2}$ represent the decay rates for the first and second-moment estimates, commonly set to $0.9$ and $0.999$, respectively. Next, we compute the bias-corrected first and second moment estimates using:
@@
Here, $\epsilon>0$ is a small constant, and $\alpha$ represents the learning rate.
Note: Algorithm 1 outlines the details of calculating $\beta$ of the proposed Wave-RVFL model. The time complexity of the proposed model is $\mathcal{O}(Itr(nmN+s)),$ with detailed information provided in Supplementary Section S.II.
@@
Input: Let $\{x_{i}\}_{i=1}^{n}$ be the input training dataset, and $Y_{i}$ be the target output. $\mathcal{C}$, $\eta$, and $\gamma$ are the trade-off parameters.
@@ 4.1 Compared Models, Datasets and Experimental Setup
To evaluate the efficacy of the proposed Wave-RVFL models, we employ $23$ benchmark datasets from the UCI [4] repository from various domains and sizes. For example, the number of samples in the “fertility” dataset is $100$, and the number of samples in the “connect_4” dataset is $67,557$. The experimental setup is discussed in Section S.III of the supplementary material.
@@ 4.2 Experimental Results on UCI Dataset
Considering the above discussion based on accuracy, rank, and statistical tests, we can conclude that the proposed Wave-RVFL model showcases superior and robust performance against the baseline models.
@@ 4.3 Robustness Evaluation of the Proposed Wave-RVFL Model on Datasets with Outliers
    “magic” Dataset: The Wave-RVFL model led the performance table in $4$ out of $4$ scenarios for the magic dataset. It recorded the highest average accuracy of $63.3254\%$, showcasing its effectiveness.
@@
These findings illustrate that the proposed Wave-RVFL model displays exceptional robustness in challenging conditions.
@@ 4.4 Robustness Evaluation of the Proposed Wave-RVFL Model on Datasets with Gaussian Noise
To evaluate the robustness of the proposed Wave-RVFL model against noise, we introduce Gaussian noise into the magic and blood datasets at levels of $5\%$, $10\%$, $15\%$, and $20\%$. This approach provided a thorough comparison of the model’s performance under adverse conditions against the RVFL, as shown in Table 3.
@@
    Overall Performance: The Wave-RVFL model’s overall average accuracy is $72.0127\%$, approximately $10\%$ higher than the RVFL model. Furthermore, it attains the lowest overall average rank of $1.25$, underscoring its superior performance among all tested models.
@@
Figure 3: Effect of hyperparameters on the proposed Wave-RVFL model.